OpenStudy (anonymous):

Please solve! 176=x(x+2)(x+11)

OpenStudy (anonymous):

Simplify the equation and write it in standard form.

OpenStudy (anonymous):

4X^2+2X-7=13X+4 4X^2+2X-13X-7-4=0 4X^2-11X-11 SOLVING USING THE QUADRATIC EQUATION X=(-B+-SQRT[B^24AC])/2A WE GET X=(11+-SQRT[121-4*4*-11])/2*4 X=(11+-SQRT[121+176])/8 X=(11+-SQRT[297)/8

OpenStudy (anonymous):

X=(11+-17.23)/8 X=(11+17.23)/8 X=28.23/8 X=3.528 SOLUTION X=(11-17.23)/8 X=-6.23/8 X=-.778 SOLUTION

OpenStudy (anonymous):

Thank you so much!